Operations Report – January 31th



Crew 274 Operations Report 1/31/2023

SOL: 2

Name of person filing report: Alexis Lojek

Non-nominal systems: Oven light, crew toilet

Notes on non-nominal systems: Oven light still needs a new bulb. We previously reported that the specifications of the oven light bulb type were emailed to mission support by crew 272 engineer. Around 1 pm the toilet started giving off a foul smell and Mission Support was contacted, as the crew suspected the toilet had an issue. Mission Support came to the habitat and assessed the situation, informing the crew engineers Lojek and O’Hara that Mission Support had to empty the toilet. There was a miscommunication regarding when the toilet requires emptying into the septic system, so Mission Support clarified what the correct answer is. We would recommend a drawing or diagram and more clear language be added to the Mission Handbook regarding the function of the habitat’s toilet. We appreciate the prompt response and clarification of toilet procedures. The toilet is now fixed and working nominally.


Spirit rover used: No

Hours: 207.5

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: 112.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: 219.5

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: No

Hours: 254.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

General notes on rovers: None of the four rovers were used on EVA. They are all charging from last night.

Summary of Hab operations: Habitat is now in nominal condition.

WATER USE: 29 gallons

Water (static tank): 416 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: During the first EVA of the day, suits 7, 8, 9, & 10 were used. During the second EVA of the day, suits 2, 3, & 7 were used. Helmets were cleaned before, between, and after EVAs. All radios were charged after mission completion and taken off the chargers after the red charging lights turned off.

Summary of GreenHab operations: GHO Hines watered plants this morning at 0846-0910 am.

WATER USE: 14 gallons used.

Heater: On

Supplemental light: 4 hours

Harvest: None

Summary of ScienceDome operations: GHO Officer Tyler Hines completed initial setup of research equipment. Experiment is being monitored.

Dual split: On.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The Musk Solar Observatory had issues in the afternoon (about 3 pm). The “black box” would not open the shutter. The Crew Astronomers included more detail in their own report. We would appreciate any advice on troubleshooting, as we plan to take a look at it tomorrow.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: The propane alarm went through its full startup cycle when power swapped over from the generator to the solar panels in the morning around 7 am. Is this something to be expected? Is this something we should be concerned about in any way (during power swap-over)? (Dr. Rupert responded to this question and stated that it was likely due to slow power turnover, will let know if it happens again). Copy on choosing a suit and using that the whole time.

Copyright © The Mars Society. All rights reserved. | Main Site